An investment banker is analyzing two companies that specialize in the production and sale of candied yams. Traditional Yams uses a labor-intensive approach, and Auto-Yams uses a mechanized system. CVP income statements for the two companies are shown below.

Traditional
Yams

Auto-Yams

Sales $409,000 $409,000
Variable costs 330,000 160,000
Contribution margin 79,000 249,000
Fixed costs 29,000 199,000
Net income $50,000 $50,000


The investment banker is interested in acquiring one of these companies. However, she is concerned about the impact that each company’s cost structure might have on its profitability.

Determine the effect on each company’s net income if sales decrease by 15% and if sales increase by 9%. Do not prepare income statements. (Round answers to 2 decimal places, e.g. 10.52. If % change is negative, enter amount with either a negative sign or parenthesis, e.g. -10.52 or (10.52).)

% Change in Net Income

Sales decrease by 15%
   Traditional Yams %
   Auto-Yams %
Sales increase by 9%
   Traditional Yams %
   Auto-Yams %
Answer #1

Sales decrease by 15%:

Traditional Yams = $79,000/$50,000 * 15%

Traditional Yams = 1.58 * 15 = 23.7%

Auto-Yams = $249,000/$50,000 * 15

Auto-Yams = 4.98 * 15 = 74.7%

Sales increase by 9%:

Traditional Yams = $79,000/$50,000 * 9

Traditional Yams = 1.58 * 9 = 14.22%

Auto-Yams = $249,000/$50,000 * 9

Auto-Yams = 4.98 * 9 = 44.82%